### v1

# HandDiffuse: Generative Controllers for Two-Hand Interactions via Diffusion Models

Abstract
Existing hands datasets are largely short-range and the interaction is weak due to the self-occlusion and self-similarity of hands, which can not yet fit the need for interacting hands motion generation. To rescue the data scarcity, we propose HandDiffuse12.5M, a novel dataset that consists of temporal sequences with strong two-hand interactions. HandDiffuse12.5M has the largest scale and richest interactions among the existing two-hand datasets. We further present a strong baseline method HandDiffuse for the controllable motion generation of interacting hands using various controllers. Specifically, we apply the diffusion model as the backbone and design two motion representations for different controllers. To reduce artifacts, we also propose Interaction Loss which explicitly quantifies the dynamic interaction process. Our HandDiffuse enables various applications with vivid two-hand interactions, i.e., motion in-betweening and trajectory control. Experiments show that our method outperforms the state-of-the-art techniques in motion generation and can also contribute to data augmentation for other datasets. Our dataset, corresponding codes, and pre-trained models will be disseminated to the community for future research towards two-hand interaction modeling.
{strip}
Figure 1: Overview: The proposed HandDiffuse12.5M benchmark dataset consists of temporal sequences with strong interaction. Based on HandDiffuse12.5M, we propose HandDiffuse, a strong baseline for the motion generation of interacting hands. HandDiffuse enables three motion generation tasks, including Unconstrained Generation, Motion In-betweening, and Trajectory Control.
1 Introduction
The tightly interacting hands play a crucial role in human emotional expression and communication, serving as a vital component for physical interaction with oneself. Controllable hand motion generation not only enhances the experience of augmented reality/virtual reality (AR/VR) but also help robotics and avatars to express their emotion in another dimension.
So far, lots of works have emerged in the field of human body motion generation [51, 59, 4, 22], but the generation of temporal hands motions in strong interaction remains blank as the lack of strong interacting hands datasets. However, the dataset acquisition is time-consuming and expensive due to the interacting hands present self-occlusion, self-similarity, and complex articulations. There are only a few datasets [53, 36, 64, 28, 37] focusing on hands with strong interactions. Most of them, i,e, InterHand2.6M [36], Two-hand 500K [64] and RenderIH [28], provides large-scale discrete frames but the available temporal sequences remain sparse. The severe lack of sufficient temporal motions makes them unsuitable for the motion generation task of two-hand interactions. Only recently, the concurrent work Re:InterHand [37] provides large-scale interaction data, yet it focuses on the motion capture of interacted hands under diverse lighting conditions.
To break the data scarcity, in this paper, we first contribute a comprehensive dataset, HandDiffuse12.5M, which focuses on capturing diverse temporal sequences with strong two-hand interactions. Our dataset consists of various modalities, including synchronous 50-view videos, 3D key points, and hand poses. It covers interacting hand motions in random and complex poses, daily communications, finger dances, and Chinese daily sign language with 250k temporal frames, resulting in roughly 12.5 million pictures. To address the issue of occlusion, which is unavoidable in hand interactions, we manually annotate the key points in 240k images. Note that our HandDiffuse12.5M dataset owns the largest scale not only in the number of temporal frames but also in the number of pictures among the existing available interacting hands datasets. Thus, it not only opens up the research for temporally consistent two-hand motion generation but also brings huge potential for future exploration in hand modeling.
Based on our novel dataset, we further provide a strong baseline algorithm named HandDiffuse, which tailors the diffusion model [21] to generate vivid interacting hands motions under diverse explicit control. Note that hand motion is an abstract form of emotional expression and ten fingers always gather together during interactions. Therefore, it is difficult to describe the motions with concrete text prompts and the control for motion generation needs to be explicit. To this end, we adopt two kinds of explicit controllers as the conditions for our diffusion models, including the discrete keyframes and the trajectories of wrists. These conditionings enable various downstream applications, ranging from hand motion inbetweening to trajectory-aided hand motion completion which connects hand motions with body motions (see Figure 1). In HandDiffuse, we design two specific motion representations for different controllers, as well as a novel Interaction Loss for the dynamic process of hand interaction, which helps to reduce the artifacts. Finally, we provide a thorough evaluation of our approach as well as a comparison with state-of-the-art motion generation methods [51, 31]. We also provide results to show that HandDiffuse can contribute to data augmentation for other datasets. Our preliminary outcomes indicate that controllable hand motion generation remains a challenging problem, and our dataset will consistently benefit further investigations of this new research direction. To summarize, our main contributions include:
  • We contribute a large-scale dataset, which provides accurate and temporally consistent tracking of human hands under diverse and strong two-hand interactions.
  • We propose a strong baseline for generating interacting hand motions from various explicit controllers, enabling vivid motion in-betweening and trajectory conditioning.
  • We introduce two motion representations as well as an Interaction Loss to significantly improve the generation results of two-hand interactions.
  • Our dataset, related codes, and pre-trained models will be made publicly available to stimulate further research.
2 Related Works
Dataset Temporal Sequence Source # Views Manually Annotation Appearance
RenderIH [28] No Synthesis - No Natural
Two-hand 500K [64] Partial IMU+Synthesis - No -
InterHand2.6M [36] Partial RGB 80 Yes Lab
Re:InterHand [37] Yes RGB 26 No Relighting
HandDiffuse12.5M(Ours) Yes RGB 50 Yes Lab
Table 1: Dataset Comparisons. We compare our HandDiffuse12.5M dataset with other existing publicly available interacting-hand datasets. Our HandDiffuse12.5M dataset contains both strong interaction and temporal sequences, and has the largest scale.
Hand Dataset.
The establishment of realistic hand datasets is a challenging and time-consuming process. However, these datasets have significantly accelerated the development of VR/AR, embodied AI and human-object interaction. Many researchers have utilized various sensors, including multiple RGB cameras, depth sensors and maker-base cameras, to collect realistic data in hand-object interaction or dual-hand interaction [36, 37, 17, 15, 5, 58, 48, 27, 3, 9, 47, 53, 40, 18, 57, 13, 11, 34].
Obtaining ground truth for specific dual-hand interaction motion from real-world captured data is challenging [36, 28]. Moon et al. [36] have attempted to combine manual and automated annotations, but manual annotations are labor-intensive. Recently, Moon et al. [37] have improved the detection accuracy of 2D keypoints and reconstructed hands to facilitate fully automated annotation schemes, but these methods are still time-consuming.
As a result, some works have established synthetic hand datasets to enhance annotation precision [64, 28, 63, 32, 10]. Among them, RenderIH [28] increases the diversity of data through re-lighting, while DART [10] provides a wide range of hand accessories to bridge the gap between synthetic and real data. However, synthetic datasets fail to balance the diversity of motion and the continuity of sequences. Some works [28, 64] sample single-frame actions for the left and right hands separately from real datasets and solve the occlusion problem between the hands through optimization. However, such optimization methods are not effective for temporal sequences. Therefore, very few papers [53, 36, 37] have collected temporal interacting hand datasets [28].
Hands Capture & Reconstruction.
The reconstruction of both hands is a significant challenge in the area of human motion capture [64]. The release of the InterHand2.6M [36] has motivated many regression-based methods [45, 1, 30, 16, 6, 8, 26, 19, 35, 35]. These regression-based methods are widely applied in VTuber and VR/AR devices due to their balance between real-time performance and accuracy. Some works [30, 38] proposed a Transformer-based network with the cross-attention between right and left hands. Recently, Zuo et al. [64] constructed a learning-based prior to capture dual-hands which achieved great results. Building on the impressive performance of diffusion models in image generation [46, 49, 21], DiffHand [29] combines hand reconstruction with diffusion models.
Human Motion generation.
In recent years, there has been a growing interest in the field of human motion generation, with many studies exploring the generation of human motions based on conditioning signals. These conditioning signals are used to guide the generation of specific types of motions, such as motion in-betweening [25, 20, 42], trajectory control [60, 43, 24, 54] and unconstrained motion synthesis [39, 55, 62]. Leading by the pioneering work of MDM [52], human motion generation models based on diffusion architecture [31, 61, 7] have been shown to gain better motion diversity and expressiveness compared to prior works based on GAN [33] or VAE [2, 41]. However, most existing works in this area have primarily focused on generating whole-body motions, neglecting the generation of dual-hand motions, which plays a crucial role in many interactive scenarios. Utilizing the capabilities of the diffusion models, we propose a novel method HandDiffuse to enable interacting hands motion generation tasks with control signals such as key frames and trajectory.
3 HandDiffuse1.2M Dataset
Figure 2: Overview of our HandDiffuse12.5M. HandDiffuse12.5M is featured with RGB images, multi-view 2D keypoints, reconstructed 3D keypoints, and MANO [44] parameters.
Figure 3: t-SNE visualization of HandDiffuse12.5M(Ours), InterHand2.6M and Re:InterHand. For InterHand2.6M, we only choose its temporal frames. The result indicates the diversity and richness of our HandDiffuse12.5M.
Figure 4: (Left) HandDiffuse Training procedure overview. Given HandDiffuse Motion Representation $\mathbf{x_{t}^{1:N}}$ of length $N$ together with a timestep embedding $t$ as input, we use a transformer Encoder to predict the final hand motion $\mathbf{\hat{x}_{0}^{1:N}}$. The proposed Interaction Loss is utilized to regulate the generated motion. (Right) Downsteam Applications overview. For Unconstrained Generation, we sample pure Gaussian noise $x_{t}$ and predict the final motion $\hat{x}_{0}$. For Motion In-betweening, we sample the final motion $\hat{x}_{0}$ given $\mathbf{x_{t}^{1:5}}$ and $\mathbf{x_{t}^{-5:N}}$. For Trajectory Control, we sample the final motion $\hat{x}_{0}$ given the root angular velocity and linear velocity.
3.1 Capture system
Existing hand motion capture solutions include IMU gloves, elastic sensor gloves, marker-based systems, and multi-view camera capture. However, due to the severe occlusion in hand interactions, marker-based systems are not suitable; IMU gloves and elastic sensor-based capture solutions theoretically address the issue of visual occlusion, but these gloves require specific hand shapes from the wearer and are difficult to calibrate in each instance of use. Therefore, we have opted for a multi-view RGB approach combined with manual annotation, similar to InterHand2.6M. As shown in Figure 3, our capture system consists of 50 mounted video cameras (Zcam E2) recording interaction sequences at a frame rate of 30 frames per second (fps). We adjust the resolution to 3840 × 2160 to allocate more pixels for capturing hand movements. To ensure uniform illumination and minimize motion blur, we utilize 35 directional LED panel lights and set the shutter speed to 1/500. We design four categories of hand interaction motions for recording: random and complex hand interactions, hand movements during daily communication, finger tutting dance, and Chinese daily sign language. Each motion sequence is controlled to last approximately one minute.
3.2 2/3D joint coordinates & MANO fitting
[…]
To address the computational challenges posed by the full MANO model, which operates with a complexity of $O(BD_{V}D_{J})$. We developed a compressed MANO model. This model reduces the computational complexity to $O(BD_{J})$ by focusing only on the joints’ positions. The efficiency is achieved by preprocessing MANO’s shape basis to formulate a ”joint shape basis”, enabling accurate hand pose depiction with fewer computations.
In the optimization phase, we aim to precisely adjust the MANO parameters ($\beta,\theta_{g},\theta_{l}$) for alignment with the captured 3D joint data. The joint positions as determined by the MANO model are represented as $J_{\text{mano}}(\beta,\theta_{g},\theta_{l})$, while the ground-truth positions are denoted as $J_{\text{gt}}$. The optimization encompasses two primary components:
$L_{\text{joint}}$ $=\lVert J_{\text{gt}}-J_{\text{mano}}(\beta,\theta_{g},\theta_{l})\rVert_{2},$
$L_{\text{pose}}$ $=\lVert\mathbf{\theta}_{l}\rVert^{2},$
$L_{\text{total}}$ $=L_{\text{joint}}+\lambda L_{\text{pose}},$
In this formulation, $\lambda>0$ is a hyperparameter that helps balance joint positioning accuracy with pose regularization.
Despite these steps, errors caused by hand occlusion are still inevitable. We manually verify each frame and annotate erroneous frames for accuracy. For each frame, we choose the six least occluded viewpoints for precise joint position annotation, totaling 240,000 pictures manually annotated. Further information about the construction of the HandDiffuse dataset and the application of the compressed MANO model for efficient parameter fitting is detailed in the Appendix.
3.3 Quantitative Evaluation for our Dataset.
Table 1 presents a comparison with two existing 3D hand datasets. Our dual-hand dataset focuses on close-hand interactions and temporal sequences. Our HandDiffuse dataset has the highest number of frames, with each sequence containing at least 200 frames. Our HandDiffuse dataset has more than 250,000 temporal frames, making it the largest known dataset of continuous temporal frames for dual-hand interaction data. We anticipate that this dataset with long sequences of close hand interactions will assist in fine-grained hands motion generation. More details are provided in the Appendix.
We evaluate the superiority of our dataset in terms of interaction data diversity and effectiveness. To this end, we compare our dataset with InterHand2.6M [36] and Re:InterHand [37], both of which also contain hand interaction motion sequences. Figure 3 illustrates that our dataset exhibits a greater variety of interaction poses compared to InterHand2.6M and Re:InterHand. This is attributed to our dataset’s explicit focus on hand interaction pose sequences for generation tasks.
4 Method
An overview of our method is shown in Figure 4. With our dataset, the goal is to synthesize two hands motion $\mathbf{x^{1:N}}$ without or with two designed controllers $c$. These two selected controllers are common in human motion generation and are also able to dictate the synthesis of two hand motions, such as initial and end frames (motion in-betweening) and the positions and rotations of both wrists (trajectory control). In contrast to single-person human motion generation, the task of generating dual-hand motion requires the network to learn not only the local motion of each hand but also the dynamic process of interaction. Therefore, we propose two specific hand motion representations for different controllers 4.1 and Interaction Loss 4.2 to model this dynamic interaction process.
Inspired by recent work, we adopt the diffusion model as the backbone which has shown remarkable generative ability in image generation and human motion generation. Diffuse is modeled as a Markov noising process and the ground truth $\mathbf{x_{0}^{1:N}}$ is diffused to $\mathbf{x_{t}^{1:N}}$ by adding t-step independent Gaussian noise $\epsilon\sim\mathcal{N}(0,I)$. This process can be formulated as
$q(\mathbf{x_{t}|x_{0}})=\sqrt{\overline{\alpha}_{t}}\mathbf{x_{0}}+(1-\overline{\alpha}_{t}),$
where $\overline{\alpha}_{t}=\prod_{s=0}^{t}\alpha_{s}$ and $\alpha_{t}=1-\beta_{t}$. $\beta_{t}$ is the cosine noise variance schedule and $q(\mathbf{x_{t}})$ is near $\mathcal{N}(0,I)$ when t is big enough. Then, $\mathbf{x_{t}}$ is sent to a denoiser conditioned on timestep $t$ and different conditions $c$ to predict $\mathbf{\hat{x}_{0}}$:
$\mathbf{\hat{x}_{0}}=\mathcal{G}(\mathbf{x_{t}},t,c).$
We use an encoder-only transformer to construct our denoiser $D$ following MDM[51] and we concatenate the left hand and right hand together as input.
4.1 Motion representation for two hands.
For different tasks, we have proposed two specific motion representation methods, which we refer to as local representation (local rep) and global representation (global rep). The local rep is inspired by HumanML3D[14], a widely used approach in human motion generation. In each frame, the pose of left hand is defined by a tuple of $(\dot{R},\dot{T},J_{local},\dot{J}_{local},\theta)$ and the right hand is $(R_{init},T_{init},\dot{R},\dot{T},J_{local},\dot{J}_{local},\theta)$, where $\dot{R}\in\mathbb{R}^{4}$ and $\dot{T}\in\mathbb{R}^{3}$ represent root angular velocity in quaternion and linear velocity respectively. $J_{local}\in\mathbb{R}^{3j}$ and $\dot{J}_{local}\in\mathbb{R}^{3j}$ denote the local joint positions and velocities in root space, with $j=21$ represents the number of joints. $\theta\in\mathbb{R}^{4i}$ represents the local joint rotations in quaternion representation, following the skeleton structure of MANO with $i=15$. Each sequence is aligned with the left wrist of the first frame as the coordinate system origin. Both left and right hands are rotated to align the left-hand R vector with the direction (1, 0, 0) which is shown in Figure 4. Therefore, the right hand has $R_{init}\in\mathbb{R}^{4}$ and $T_{init}\in\mathbb{R}^{3}$.
On the other hand, the global rep is defined as a tuple of ($J_{global}$, $\dot{J}_{global}$, $\theta$) for both hands, where $J_{global}$ represents the positions of hand joints in the world coordinate system and two hands are normalized by employing the same approach as local rep.
Figure 5: Module of contact potential loss in Interaction Loss. The distance between joints has direction. Despite the approximate similarity in length of $\overrightarrow{d_{1}}$ and $\overrightarrow{d_{2}}$, their directions are opposite. Therefore, it is necessary to consider both the length and direction simultaneously to eliminate ambiguity.
Figure 6: The Downstream Application of HandDiffuse. The generation results of downstream applications of HandDiffuse. The generated dual-hand motion exhibits natural and realistic hand interaction.
During the experimental process, we observed that the local rep performed better in the task of trajectory control, where the global rotation and translation of the hand are given. The global rep achieved better results in the unconditional generation and motion in-betweening tasks. In Section 5.3, a detailed comparison between the two representations is provided.
4.2 Interaction Loss.
The first part of the Interaction Loss is the dual-hand joint contact potential loss. In each frame, we first calculate a $25\times 25$ distance matrix $\mathbf{D}$ between the 21 left-hand joints and 21 right-hand joints with 4 more joints sampled in each palm. Since the interaction between the hands involves fine-grained movements, we should focus on situations where the finger joints are very close to each other. Inspired by the concept of Contact Potential Field[56], we further represent the distances using the elastic potential energy of a spring model and it can be described as
$\mathbf{P}otential=1/2\mathbf{K}\ relu(\mathbf{\tau}-\|\mathbf{D}_{pred}\|_{2})^{2},$ (1)
where $\mathbf{K}$ represents the coefficient of elasticity and $\mathbf{\tau}$ denotes the distance threshold. The closer the distance between the finger joints, the larger the potential energy obtained, and the distance will be filtered out when it is over the threshold. Moreover, by considering the direction of the distance, we can avoid ambiguity caused by penetration, and the final contact potential loss is described as
$\mathbf{Loss}_{potential}=|\mathbf{P}_{pred}-\mathbf{P}_{gt}|(1+\overrightarrow{\mathbf{D}_{pred}}\times\overrightarrow{\mathbf{D}_{gt}}),$ (2)
where (×)($\times$)( × ) means cross product. More is shown in Figure 5.
Figure 7: Qualitative comparison for different methods. For Unconstrained Generation, MDM[52] tend to generate motion with hands shifting away from each other, while InterGen[31] would generate motions with a mean pose. In terms of Trajectory Control, MDM yields satisfactory results, whereas InterGen generates motion without incorporating interaction.
During our testing process, we observed that MANO has strict requirements for the proportional lengths of hand bones. Therefore, the second part of the Interaction Loss is the Shape Loss. We ensure the rationality of hand shape by calculating the differences in bone lengths between the left and right hands, as well as the differences between the predicted shape and the ground truth (GT). This is described as follows:
$\mathbf{Loss}_{shape}=|\mathbf{BL}_{left}-\mathbf{BL}_{right}|+|\mathbf{BL}_{pred}-\mathbf{BL}_{GT}|,$ (3)
where $BL$ denotes the Euclidean distance of the bones.
4.3 Downstream Application
For the task of unconstrained generation and motion in-betweening, the key challenge lies in constructing the global information of the dynamic interaction process between the two hands. Therefore, we choose to use the global rep. In motion in-betweening task, we provide the ground truth for the previous five frames $\mathbf{x_{0}^{1:5}}$ and the last five frames $\mathbf{x_{0}^{-5:-1}}$ for the diffused motion $\mathbf{x_{t}^{1:N}}$.
For the task of trajectory control, the goal is to generate the reasonable local motion given the global rotation and translation of both hands. In this case, using the local rep allows the network to focus on the local motion during the interaction process. Furthermore, we provide $\dot{R}^{1:N}$ and $\dot{T}^{1:N}$ for the diffuse $\mathbf{x_{t}^{1:N}}$.
5 Experiment
In this section, we conduct several experiments to evaluate our dataset and method for dual-hand motion generation tasks. We particularly evaluate (1) the comparison of our method against previous state-of-the-art approaches, (2) the ability of our method to augment existing datasets, and (3) the effectiveness of our hand motion representation and Interaction Loss. The above evaluations are conducted specifically on the three tasks mentioned in Section 4.3. The models have been trained with $T=1000$ noising steps and a cosine noise schedule. All of them have been trained on a single NVIDIA GeForce RTX 3080 GPU for about two days.
Methods Unconstrained In-betweening Trajectory Control
[…]
Real 0.050 12.034
MDM[52] 0.526 10.366 0.616 10.258 0.205 11.265
InterGen[31] 0.251 4.928 0.269 5.385 0.402 9.102
Ours 0.176 11.786 0.282 11.533 0.182 11.523
Table 2: Quantitative comparisons of different methods. We compare the FID and Diversity for different methods. ’normal-$\downarrow$’ indicates that results are better with lower metrics. Bold means the best result.
5.1 Comparisons of different methods
To the best of our knowledge, our proposed method is the first to specifically target hand-only interaction motion generation. To provide a comprehensive evaluation, we compare our approach with two recent state-of-the-art human motion generation methods, namely MDM [52] and InterGen [31]. We modify these methods by adapting their motion representation to enable hand motion generation. As depicted in Figure 7, both MDM and InterGen exhibit limitations when generating hand motion. They tend to produce motion sequences with severe joint penetration or unnatural interaction since they were not originally designed for dual-hand motion generation. In contrast, our proposed method showcases superior performance as demonstrated by the lowest Fréchet Inception Distance (FID) values, as shown in Table 2. The way we adopted to calculate FID is introduced in the Appendix.
Figure 8: Cross data Evaluation. Motion In-betweening task result on InterHand2.6M[36]. We sample two frames in InterHand2.6M as the start frame and end frame and generate the in-betweening motion. Our HandDiffse can help augment the existing datasets.
5.2 Cross Datasets Evaluation
To evaluate our method cross datasets, we conduct experiments on two large-scale dual-hand interaction motion datasets, InterHand2.6M[36] and Re:InterHand [37] by sampling control signal from these two datasets. As shown in Figure 8, given the first frame and last frame from InterHand2.6M, our model is able to generate diverse motions. Furthermore, Table 3 shows that after performing In-betweening and Trajectory Control tasks on the datasets, the generated data can increase the diversity. This demonstrates the effectiveness of our approach and HandDiffuse can contribute to data augmentation for other datasets. More results and the details of experiments are introduced in the Appendix.
Datasets InterHand Re:InterHand
[…]
Real 0.113 8.478 0.034 10.583
In-betweening 0.194 8.171 0.208 11.200
Root Control 0.142 10.146 0.236 12.979
Table 3: Cross Dataset Evaluation.The FID and Diversity both increase, indicating that we get a larger dataset from the original dataset by performing data generation tasks.
5.3 Ablation Study
As we propose two kinds of hand motion representations in section 4.1, here we demonstrate the effectiveness of the representations in different tasks by conducting qualitative and quantitative comparisons. Figure 9 illustrates the limitations of the model without Shape Loss and Interaction Loss. By incorporating the Interaction Loss, we mitigate the issue of joint penetration, resulting in improved motion generation quality. To provide a quantitative analysis, we employ FID and the penetration metric sdf loss(SDF)[23] for evaluation. SDF represents the Signed Distance Field, which could be defined as:
$\phi(x,y,z)=-\min(SDF(x,y,z),0),$ (4)
where $(x,y,z)$ denotes the position of MANO vertex. According to the definition, when the hand joints penetrate with each other, $\phi$ would take the positive value while zero otherwise. Using the formulation, we can measure the severity of penetration during the whole motion. Table 4 demonstrates that our model achieves the highest performance when using the proper motion representation and Interaction Loss.
Tasks Unconstrained In-betweening Trajectory Control
[…]
Real 0.050 1.146 - -
w Global rep. 0.168 1.575 0.282 0.415
w Local rep. 0.436 1.988 0.572 0.182
w/o Interaction Loss 0.176 1.652 0.318 0.216
Complete model 0.168 1.575 0.282 0.182
Table 4: Quantitative comparisons of different hand motion representations and w/o Interaction Loss. SDF indicates the sdf loss for penetration of joints during the whole motion. Note that the mean of sdf loss is only calculated on frames with penetraion.
6 Conclusion
Figure 9: Ablation Study. In the absence of Shape Loss, the model generates abnormal hand shapes. Similarly, the absence of Interaction Loss leads to hand interaction motion with severe penetration.
We present HandDiffuse12.5M, the largest interacting hands dataset with strong interaction and temporal sequences, to open up the research direction of interacting hands motion generation. HandDiffuse12.5M consists of various modalities, including synchronous 50-view videos, 3D keypoints and hand poses. Based on HandDiffuse12.5M, we propose a strong baseline method HandDiffuse which is the first interacting hands motion generation approach along with two specific controllers: key frames control and trajectory control. Our evaluation in multi-stage demonstrates the benefit of our approach againist others and HandDiffuse fills the blank in human motion generation.
Appendix
7 Experiment Details
7.1 Frechet Inception Distance (FID)
The Frechet Inception Distance (FID) is a metric used to quantify the dissimilarity between the distributions of generated and ground truth data in latent space. In our experiments, we employ FID as a crucial indicator to measure the authenticity and plausibility of our generated data. And the formula is
$FID=|\mu_{gt}-\mu_{gen}|^{2}+\text{Tr}(C_{gt}+C_{gen}-2\sqrt{C_{gt}C_{gen}}),$
where $\mu_{gt}$ and $\mu_{gen}$ are the mean feature vectors of ground truth and generated motions, respectively; $C_{gt}$ and $C_{gen}$ are the covariance feature matrices of real and generated motions, respectively; Tr()($\cdot$)( ⋅ ) denotes the trace of the matrix. In the previous works[51], the feature is extracted by a Inception network which is a classifier. However, in the interacting hands datasets, it is challenging to label these complex interactive movements, making it difficult to build a classifier. Therefore, taking inspiration from MotionCLIP[50], we propose a self-supervised encoder-decoder architecture for the reconstruction network, as depicted in Figure 10.
More specifically, we concatenate a $z_{tk}$ vector of equal dimensions to the original motion representation $x^{1:T}$. This concatenated sequence $x^{1:T+1}$ is then fed into a transformer encoder. From the encoder’s output, we select the corresponding position $z_{x}$ as the feature representation for the input motion segment. This obtained feature $z_{x}$ serves as the memory for the decoder. Additionally, we use a T-length zero vector that undergoes positional encoding as the target for reconstructing the motion in the decoder. The reconstruction loss is computed using the L2 norm.
During experimentation, we observed that using the global representation of the motion improves the convergence of the network. Therefore, in subsequent experiments, we convert the motion to a global representation before calculating the motion’s feature.
Figure 10: Feature Extractor for FID. Inspired by MotionCLIP[50], we train a transformer with encoder-decoder architecture by using self-supervised strategy.
7.2 Diversity
Diversity measures the variability in the resulting motion distribution. The way we followed to calculate Diversity is as same as MDM[51]. We randomly generate 1000 samples and select 200 of them for calculating.
8 HandDiffuse12.5M
Figure 11 has shown the diversity and complexity of our dataset HandDiffuse12.5M. Figure 12 has shown our manual annotation tool.
Figure 11: Overview of HandDiffuse12.5M. Each image represents one frame of hand motion from a different sequence.
Figure 12: Annotation Tool. Our specially designed tool for hand joint annotation of severely occluded frames.
9 More Results
Figure 13 has shown the results of trajectory control task. In each sequence, we adopt the wrists’ rotation and translation from InterHand2.6M[36] as our control signal and we can generate different motion based on this trajectory.
Figure 13: Results of Trajectory Control by Sampling from InterHand2.6M. Note that only Root orientation and position is given, the local motions for hand joints are generated from our model.

### v2

@@
 Existing hands datasets are largely short-range and the interaction is weak due to the self-occlusion and self-similarity of hands, which can not yet fit the need for interacting hands motion generation.
-To rescue the data scarcity, we propose HandDiffuse12.5M, a novel dataset that consists of temporal sequences with strong two-hand interactions.
+To rescue the data scarcity, we propose HandDiffuse12.5M, a novel and real dataset that consists of temporal sequences with strong two-hand interactions.
 HandDiffuse12.5M has the largest scale and richest interactions among the existing two-hand datasets.
@@
 To reduce artifacts, we also propose Interaction Loss which explicitly quantifies the dynamic interaction process.
-Our HandDiffuse enables various applications with vivid two-hand interactions, i.e., motion in-betweening and trajectory control.
-Experiments show that our method outperforms the state-of-the-art techniques in motion generation and can also contribute to data augmentation for other datasets.
-Our dataset, corresponding codes, and pre-trained models will be disseminated to the community for future research towards two-hand interaction modeling.
-{strip}
-Figure 1: Overview: The proposed HandDiffuse12.5M benchmark dataset consists of temporal sequences with strong interaction.
-Based on HandDiffuse12.5M, we propose HandDiffuse, a strong baseline for the motion generation of interacting hands.
-HandDiffuse enables three motion generation tasks, including Unconstrained Generation, Motion In-betweening, and Trajectory Control.
-1 Introduction
+Our HandDiffuse enables various applications, i.e., motion in-betweening and trajectory controled generation.
+Experiments show that our method outperforms the state-of-the-art techniques in motion generation.
+The vivid two-hand motions generated by our method can also construct synthetic datasets and enhances the accuracy of existing hand motion capture algorithms.
+Datasetshttps://handdiffuse.github.io/
+Introduction
 The tightly interacting hands play a crucial role in human emotional expression and communication, serving as a vital component for physical interaction with oneself.
 Controllable hand motion generation not only enhances the experience of augmented reality/virtual reality (AR/VR) but also help robotics and avatars to express their emotion in another dimension.
-So far, lots of works have emerged in the field of human body motion generation [51, 59, 4, 22], but the generation of temporal hands motions in strong interaction remains blank as the lack of strong interacting hands datasets.
+So far, lots of works have emerged in the field of human body motion generation (Tevet et al.
+2022b; Zhang et al.
+2022a; Chen et al.
+2023; Jiang et al.
+2023), but the generation of temporal hands motions in strong interaction remains blank as the lack of strong interacting hands datasets.
 However, the dataset acquisition is time-consuming and expensive due to the interacting hands present self-occlusion, self-similarity, and complex articulations.
-There are only a few datasets [53, 36, 64, 28, 37] focusing on hands with strong interactions.
-Most of them, i,e, InterHand2.6M [36], Two-hand 500K [64] and RenderIH [28], provides large-scale discrete frames but the available temporal sequences remain sparse.
+There are only a few datasets (Tzionas et al.
+2016; Moon et al.
+2020; Zuo et al.
+2023; Li et al.
+2023a; Moon et al.
+2023) focusing on hands with strong interactions.
+Most of them, i,e, InterHand2.6M (Moon et al.
+2020), Two-hand 500K (Zuo et al.
+2023) and RenderIH (Li et al.
+2023a), provides large-scale discrete frames but the available temporal sequences remain sparse.
 The severe lack of sufficient temporal motions makes them unsuitable for the motion generation task of two-hand interactions.
-Only recently, the concurrent work Re:InterHand [37] provides large-scale interaction data, yet it focuses on the motion capture of interacted hands under diverse lighting conditions.
-To break the data scarcity, in this paper, we first contribute a comprehensive dataset, HandDiffuse12.5M, which focuses on capturing diverse temporal sequences with strong two-hand interactions.
+Only recently, the concurrent work Re:InterHand (Moon et al.
+2023) provides large-scale interaction data, yet it focuses on the motion capture of interacted hands under diverse lighting conditions.
+Figure 1: Overview: The proposed HandDiffuse12.5M benchmark dataset consists of temporal sequences with strong interaction.
+Based on it, we propose HandDiffuse, a strong baseline for the motion generation of interacting hands.
+To break the data scarcity, in this paper, we first contribute a real dataset, HandDiffuse12.5M, which focuses on capturing diverse temporal sequences with strong two-hand interactions.
 Our dataset consists of various modalities, including synchronous 50-view videos, 3D key points, and hand poses.
-It covers interacting hand motions in random and complex poses, daily communications, finger dances, and Chinese daily sign language with 250k temporal frames, resulting in roughly 12.5 million pictures.
+It covers interacting hand motions in random and complex poses, daily communications, finger dances, and Chinese daily sign language with 250K temporal frames($\approx$ 2 hours), resulting in roughly 12.5M pictures.
 To address the issue of occlusion, which is unavoidable in hand interactions, we manually annotate the key points in 240K images.
-Note that our HandDiffuse12.5M dataset owns the largest scale not only in the number of temporal frames but also in the number of pictures among the existing available interacting hands datasets.
+Note that our HandDiffuse12.5M dataset owns the largest scale not only in the duration but also in the number of pictures among the existing available interacting hands datasets.
 Thus, it not only opens up the research for temporally consistent two-hand motion generation but also brings huge potential for future exploration in hand modeling.
-Based on our novel dataset, we further provide a strong baseline algorithm named HandDiffuse, which tailors the diffusion model [21] to generate vivid interacting hands motions under diverse explicit control.
+Based on our novel dataset, we further provide a strong baseline algorithm named HandDiffuse, which tailors the diffusion model (Ho, Jain, and Abbeel 2020) to generate vivid interacting hands motions under diverse explicit control.
 Note that hand motion is an abstract form of emotional expression and ten fingers always gather together during interactions.
 Therefore, it is difficult to describe the motions with concrete text prompts and the control for motion generation needs to be explicit.
-To this end, we adopt two kinds of explicit controllers as the conditions for our diffusion models, including the discrete keyframes and the trajectories of wrists.
-These conditionings enable various downstream applications, ranging from hand motion inbetweening to trajectory-aided hand motion completion which connects hand motions with body motions (see Figure 1).
-In HandDiffuse, we design two specific motion representations for different controllers, as well as a novel Interaction Loss for the dynamic process of hand interaction, which helps to reduce the artifacts.
-Finally, we provide a thorough evaluation of our approach as well as a comparison with state-of-the-art motion generation methods [51, 31].
+To this end, we adopt several explicit controllers as the conditions for our diffusion models, including the discrete keyframes, the trajectories of wrists and 2D keypoints.
+These conditionings enable various downstream applications, ranging from hand motion in-betweening to trajectory-aided hand motion completion which connects hand motions with body motions (see Figure1), and can also assist in hand motion reconstruction.
+In HandDiffuse, we propose Interacting Hands DDIM which is inspired by modular design and come up with two specific motion representations for different controllers, as well as a novel Interaction Loss for the dynamic process of hand interaction, which helps to reduce the artifacts.
+Finally, we provide a thorough evaluation of our approach as well as a comparison with state-of-the-art motion generation methods (Tevet et al.
+2022b; Liang et al.
+2023; Lee et al.
+2024; Müller et al.
+2024).
 We also provide results to show that HandDiffuse can contribute to data augmentation for other datasets.
@@
 To summarize, our main contributions include:
-  • We contribute a large-scale dataset, which provides accurate and temporally consistent tracking of human hands under diverse and strong two-hand interactions.
-  • We propose a strong baseline for generating interacting hand motions from various explicit controllers, enabling vivid motion in-betweening and trajectory conditioning.
-  • We introduce two motion representations as well as an Interaction Loss to significantly improve the generation results of two-hand interactions.
-  • Our dataset, related codes, and pre-trained models will be made publicly available to stimulate further research.
-2 Related Works
+  • We contribute a large-scale real dataset named HandDiffuse12.5M, which provides accurate and temporally consistent tracking of human hands under diverse and strong two-hand interactions.
+  • We are the first to specifically target on hand-only interaction motion generation and we propose a strong baseline for generating interacting hand motions from various explicit controllers, enabling vivid motion in-betweening, trajectory conditioning and hands reconstruction.
+  • We introduce Interacting Hands DDIM, two motion representations as well as an Interaction Loss to significantly improve the generation results of two-hand interactions.
+The generated motions can contribute to data augmentation for other datasets.
+  • Our dataset, related codes will be made publicly available to stimulate further research.
+Related Works
 Dataset Temporal Sequence Source # Views Manually Annotation Appearance
-RenderIH [28] No Synthesis - No Natural
-Two-hand 500K [64] Partial IMU+Synthesis - No -
-InterHand2.6M [36] Partial RGB 80 Yes Lab
-Re:InterHand [37] Yes RGB 26 No Relighting
+RenderIH (Li et al. 2023a) No Synthesis - No Natural
+Two-hand 500K (Zuo et al. 2023) Partial IMU+Synthesis - No -
+InterHand2.6M (Moon et al. 2020) Partial RGB 80 Yes Lab
+Re:InterHand (Moon et al. 2023) Yes RGB 26 No Relighting
 HandDiffuse12.5M(Ours) Yes RGB 50 Yes Lab
@@
 Hand Dataset.
-The establishment of realistic hand datasets is a challenging and time-consuming process.
-However, these datasets have significantly accelerated the development of VR/AR, embodied AI and human-object interaction.
-Many researchers have utilized various sensors, including multiple RGB cameras, depth sensors and maker-base cameras, to collect realistic data in hand-object interaction or dual-hand interaction [36, 37, 17, 15, 5, 58, 48, 27, 3, 9, 47, 53, 40, 18, 57, 13, 11, 34].
-Obtaining ground truth for specific dual-hand interaction motion from real-world captured data is challenging [36, 28].
+Although hand datasets have significantly accelerated the development of VR/AR, embodied AI and human-object interaction.
+Obtaining ground truth for specific dual-hand interaction motion from real-world captured data is challenging (Moon et al.
+2020; Li et al.
+2023a).
 Moon et al.
-[36] have attempted to combine manual and automated annotations, but manual annotations are labor-intensive.
+(Moon et al.
+2020) have attempted to combine manual and automated annotations, but manual annotations are labor-intensive.
 Recently, Moon et al.
-[37] have improved the detection accuracy of 2D keypoints and reconstructed hands to facilitate fully automated annotation schemes, but these methods are still time-consuming.
-As a result, some works have established synthetic hand datasets to enhance annotation precision [64, 28, 63, 32, 10].
-Among them, RenderIH [28] increases the diversity of data through re-lighting, while DART [10] provides a wide range of hand accessories to bridge the gap between synthetic and real data.
+(Moon et al.
+2023) have improved the detection accuracy of 2D keypoints and reconstructed hands to facilitate fully automated annotation schemes, but these methods are still time-consuming.
+As a result, some works have established synthetic hand datasets to enhance annotation precision (Zuo et al.
+2023; Li et al.
+2023a; Zimmermann and Brox 2017; Lin, Wilhelm, and Martinez 2021; Gao et al.
+2022).
+Among them, RenderIH (Li et al.
+2023a) increases the diversity of data through re-lighting, while DART (Gao et al.
+2022) provides a wide range of hand accessories to bridge the gap between synthetic and real data.
 However, synthetic datasets fail to balance the diversity of motion and the continuity of sequences.
-Some works [28, 64] sample single-frame actions for the left and right hands separately from real datasets and solve the occlusion problem between the hands through optimization.
+Some works (Li et al.
+2023a; Zuo et al.
+2023) sample single-frame actions for the left and right hands separately from real datasets and solve the occlusion problem between the hands through optimization.
 However, such optimization methods are not effective for temporal sequences.
-Therefore, very few papers [53, 36, 37] have collected temporal interacting hand datasets [28].
+Therefore, very few papers (Tzionas et al.
+2016; Moon et al.
+2020, 2023) have collected temporal interacting hand datasets (Li et al.
+2023a).
 Hands Capture & Reconstruction.
-The reconstruction of both hands is a significant challenge in the area of human motion capture [64].
-The release of the InterHand2.6M [36] has motivated many regression-based methods [45, 1, 30, 16, 6, 8, 26, 19, 35, 35].
-These regression-based methods are widely applied in VTuber and VR/AR devices due to their balance between real-time performance and accuracy.
-Some works [30, 38] proposed a Transformer-based network with the cross-attention between right and left hands.
+The reconstruction of both hands is a significant challenge in the area of human motion capture (Zuo et al.
+2023).
+The release of the InterHand2.6M (Moon et al.
+2020) has motivated many regression-based methods (Rong et al.
+2021; Baowen et al.
+2021; Li et al.
+2022; Hampali et al.
+2022; Di and Yu 2022; Fan et al.
+2021; Kim, In Kim, and Baek 2021; Hao et al.
+2022; Moon 2023).
+These regression-based methods significantly boost the development of VTuber and VR/AR devices due to their balance between real-time performance and accuracy.
+Some works (Li et al.
+2022; Park et al.
+2023) proposed a Transformer-based network with the cross-attention between right and left hands.
 Recently, Zuo et al.
-[64] constructed a learning-based prior to capture dual-hands which achieved great results.
-Building on the impressive performance of diffusion models in image generation [46, 49, 21], DiffHand [29] combines hand reconstruction with diffusion models.
+(Zuo et al.
+2023) constructed a learning-based prior to capture dual-hands which achieved great results.
+Building on the impressive performance of diffusion models in image generation (Ronneberger, Fischer, and Brox 2015; Song, Meng, and Ermon 2020; Ho, Jain, and Abbeel 2020), DiffHand (Li et al.
+2023b) combines hand reconstruction with diffusion models.
+The existing hands capture algorithms can also get enhanced by synthetic data.
+RenderIH(Li et al.
+2023a) and DART(Gao et al.
+2022) have shown that the synthetic hand datasets can improve the capture accuracy.
 Human Motion generation.
 In recent years, there has been a growing interest in the field of human motion generation, with many studies exploring the generation of human motions based on conditioning signals.
-These conditioning signals are used to guide the generation of specific types of motions, such as motion in-betweening [25, 20, 42], trajectory control [60, 43, 24, 54] and unconstrained motion synthesis [39, 55, 62].
-Leading by the pioneering work of MDM [52], human motion generation models based on diffusion architecture [31, 61, 7] have been shown to gain better motion diversity and expressiveness compared to prior works based on GAN [33] or VAE [2, 41].
-However, most existing works in this area have primarily focused on generating whole-body motions, neglecting the generation of dual-hand motions, which plays a crucial role in many interactive scenarios.
+These conditioning signals are used to guide the generation of specific types of motions, such as motion in-betweening (Kaufmann et al.
+2020; Harvey et al.
+2020; Qin, Zheng, and Zhou 2022), trajectory control (Zhang et al.
+2021; Rempe et al.
+2023; Karunratanakul et al.
+2023; Xie et al.
+2023) and unconstrained motion synthesis (Pavlakos et al.
+2019; Yan et al.
+2019; Zhao, Su, and Ji 2020).
+Leading by the pioneering work of MDM (Tevet et al.
+2023), human motion generation models based on diffusion architecture (Liang et al.
+2023; Zhang et al.
+2022b; Du et al.
+2023) have been shown to gain better motion diversity and expressiveness compared to prior works based on GAN (Lin and Amer 2018) or VAE (Cai et al.
+2021; Petrovich, Black, and Varol 2021).
+However, most existing works in this area have primarily focused on generating whole-body motions, only few works(Lee et al.
+2024; Zhang et al.
+2024) try to generate hands’ motion in recent.
 Utilizing the capabilities of the diffusion models, we propose a novel method HandDiffuse to enable interacting hands motion generation tasks with control signals such as key frames and trajectory.
-3 HandDiffuse1.2M Dataset
-Figure 2: Overview of our HandDiffuse12.5M. HandDiffuse12.5M is featured with RGB images, multi-view 2D keypoints, reconstructed 3D keypoints, and MANO [44] parameters.
+HandDiffuse12.5M Dataset
+Capture system
+Existing hand motion capture solutions include IMU gloves, elastic sensor gloves, marker-based systems, and multi-view camera capture.
+However, due to the severe occlusion in hand interactions, marker-based systems are not suitable; IMU gloves and elastic sensor-based capture solutions theoretically address the issue of visual occlusion, but the IMU gloves are strict to calibration and it is easy to accidentally touch the sensor during the hands interaction; the elastic sensor-based capture solutions lacks the restoration of all degrees of freedom for finger joints.
+Therefore, we have opted for a multi-view RGB approach combined with manual annotation, similar to InterHand2.6M.
+Our capture system consists of 50 mounted video cameras recording interaction sequences at a frame rate of 30-60 frames per second (fps).
+We adjust the resolution to 3840 × 2160 to allocate more pixels for capturing hand movements.
+We design four categories of hand interaction motions for recording: random but complex hand interactions, daily communication, finger tutting dance, and sign language.
+Each motion sequence is controlled to last approximately one minute limited by the hardware.
+More detailed description about the capture system has been shown in Appendix.
+2/3D joint coordinates & MANO fitting
+Upon securing multi-view video sequences, we utilize DWPose(Yang et al.
+2023) which is the SOTA 2D keypoints detector to extract the hands 2D keypoints.
+The subsequent triangulation of these 2D keypoints yields 3D keypoints in world coordinate, which are then integrated into 3D pose by optimizing MANO (Romero, Tzionas, and Black 2017) which is differentiable.
+The frames with failed fitting will be selected out and manually re-annotated with 2D key points on the images.
+Figure2 has shown some MANO reprojection results which contains strong and various hands interaction.
+Due to space limitation, a more detailed procedure for annotating the dataset and the approach we used to fit MANO is illustrated in the Appendix.
+Figure 2: The capture system and reprojectoin of MANO. The proposed HandDiffuse12.5M benchmark dataset consists of strong and various interaction with accurate annotation.
+Quantitative Evaluation for our Dataset.
+Our HandDiffuse12.5M dataset focuses on temporal sequences which has more than 250K temporal frames.
+This makes it the largest dataset among all of the interacting hands datasets as presented in Table 1.
+As shown in Figure 3, each sequence has around 1000 to 4000 temporal frames and the average is 2955 temporal frames per motion sequence, with the median of 3630 frames.
+Figure 3: The distribution of HandDiffuse12.5M’s temporal frames.
 Figure 4: t-SNE visualization of HandDiffuse12.5M(Ours), InterHand2.6M and Re:InterHand.
 For InterHand2.6M, we only choose its temporal frames.
-The result indicates the diversity and richness of our HandDiffuse12.5M.
-Figure 4: (Left) HandDiffuse Training procedure overview.
-Given HandDiffuse Motion Representation $\mathbf{x_{t}^{1:N}}$ of length $N$ together with a timestep embedding $t$ as input, we use a transformer Encoder to predict the final hand motion $\mathbf{\hat{x}_{0}^{1:N}}$.
-The proposed Interaction Loss is utilized to regulate the generated motion.
-(Right) Downsteam Applications overview.
-For Unconstrained Generation, we sample pure Gaussian noise $x_{t}$ and predict the final motion $\hat{x}_{0}$.
-For Motion In-betweening, we sample the final motion $\hat{x}_{0}$ given $\mathbf{x_{t}^{1:5}}$ and $\mathbf{x_{t}^{-5:N}}$.
-For Trajectory Control, we sample the final motion $\hat{x}_{0}$ given the root angular velocity and linear velocity.
-3.1 Capture system
-Existing hand motion capture solutions include IMU gloves, elastic sensor gloves, marker-based systems, and multi-view camera capture.
-However, due to the severe occlusion in hand interactions, marker-based systems are not suitable; IMU gloves and elastic sensor-based capture solutions theoretically address the issue of visual occlusion, but these gloves require specific hand shapes from the wearer and are difficult to calibrate in each instance of use.
-Therefore, we have opted for a multi-view RGB approach combined with manual annotation, similar to InterHand2.6M.
-As shown in Figure 3, our capture system consists of 50 mounted video cameras (Zcam E2) recording interaction sequences at a frame rate of 30 frames per second (fps).
-We adjust the resolution to 3840 × 2160 to allocate more pixels for capturing hand movements.
-To ensure uniform illumination and minimize motion blur, we utilize 35 directional LED panel lights and set the shutter speed to 1/500.
-We design four categories of hand interaction motions for recording: random and complex hand interactions, hand movements during daily communication, finger tutting dance, and Chinese daily sign language.
-Each motion sequence is controlled to last approximately one minute.
-3.2 2/3D joint coordinates & MANO fitting
-[…]
-To address the computational challenges posed by the full MANO model, which operates with a complexity of $O(BD_{V}D_{J})$.
-We developed a compressed MANO model.
-This model reduces the computational complexity to $O(BD_{J})$ by focusing only on the joints’ positions.
-The efficiency is achieved by preprocessing MANO’s shape basis to formulate a ”joint shape basis”, enabling accurate hand pose depiction with fewer computations.
-In the optimization phase, we aim to precisely adjust the MANO parameters ($\beta,\theta_{g},\theta_{l}$) for alignment with the captured 3D joint data.
-The joint positions as determined by the MANO model are represented as $J_{\text{mano}}(\beta,\theta_{g},\theta_{l})$, while the ground-truth positions are denoted as $J_{\text{gt}}$.
-The optimization encompasses two primary components:
-$L_{\text{joint}}$ $=\lVert J_{\text{gt}}-J_{\text{mano}}(\beta,\theta_{g},\theta_{l})\rVert_{2},$
-$L_{\text{pose}}$ $=\lVert\mathbf{\theta}_{l}\rVert^{2},$
-$L_{\text{total}}$ $=L_{\text{joint}}+\lambda L_{\text{pose}},$
-In this formulation, $\lambda>0$ is a hyperparameter that helps balance joint positioning accuracy with pose regularization.
-Despite these steps, errors caused by hand occlusion are still inevitable.
-We manually verify each frame and annotate erroneous frames for accuracy.
-For each frame, we choose the six least occluded viewpoints for precise joint position annotation, totaling 240,000 pictures manually annotated.
-Further information about the construction of the HandDiffuse dataset and the application of the compressed MANO model for efficient parameter fitting is detailed in the Appendix.
-3.3 Quantitative Evaluation for our Dataset.
-Table 1 presents a comparison with two existing 3D hand datasets.
-Our dual-hand dataset focuses on close-hand interactions and temporal sequences.
-Our HandDiffuse dataset has the highest number of frames, with each sequence containing at least 200 frames.
-Our HandDiffuse dataset has more than 250,000 temporal frames, making it the largest known dataset of continuous temporal frames for dual-hand interaction data.
-We anticipate that this dataset with long sequences of close hand interactions will assist in fine-grained hands motion generation.
-More details are provided in the Appendix.
-We evaluate the superiority of our dataset in terms of interaction data diversity and effectiveness.
-To this end, we compare our dataset with InterHand2.6M [36] and Re:InterHand [37], both of which also contain hand interaction motion sequences.
+The result indicates the diversity and richness of HandDiffuse12.5M.
+Our dataset’s average shortest distance between two hands’ joints is 3.61cm, superior to InterHand2.6M’s 4.04 cm and Re:InterHand’s 3.84 cm.
+We classify a sample as “contacting” if the shortest distance between two hands’ vertices is under 3 mm.
+With a contacting hands ratio of 57.3%, our dataset exceeds InterHand2.6M’s 44.74% and Re:InterHand’s 52.7%.
+We randomly selected 5000 images from the dataset to manually annotate the 2D keypoints.
+The error between the annotation results and the reprojection of 3D keypoints is 11.3 pixels per keypoint in a 3840×2160 image space.
+The MANO fitting error is 4.3mm per joint.
+We also compare our dataset with InterHand2.6M (Moon et al.
+2020) and Re:InterHand (Moon et al.
+2023), both of which also contain hand interaction motion sequences, by visulizing the t-SNE results shown in Figure 4.
 Figure 4 illustrates that our dataset exhibits a greater variety of interaction poses compared to InterHand2.6M and Re:InterHand.
-This is attributed to our dataset’s explicit focus on hand interaction pose sequences for generation tasks.
-4 Method
-An overview of our method is shown in Figure 4.
-With our dataset, the goal is to synthesize two hands motion $\mathbf{x^{1:N}}$ without or with two designed controllers $c$.
-These two selected controllers are common in human motion generation and are also able to dictate the synthesis of two hand motions, such as initial and end frames (motion in-betweening) and the positions and rotations of both wrists (trajectory control).
-In contrast to single-person human motion generation, the task of generating dual-hand motion requires the network to learn not only the local motion of each hand but also the dynamic process of interaction.
-Therefore, we propose two specific hand motion representations for different controllers 4.1 and Interaction Loss 4.2 to model this dynamic interaction process.
+This is attributed to our dataset’s explicit focus on hand interaction sequences for generation tasks.
+In summary, HandDiffuse12.5M comprises 12.5 million images and 250K temporal frames, making it the largest two-hand dataset to date.
+It provides a variety of strongly interactive motion and accurate annotations.
+Method
+Figure 5: (Left) Overview of HandDiffuse We first generate the motions for each hand separately by training Single Hand Denoiser which only focuses on local poses.
+The generated two single hands’ local pose are concatenated with global information(noise) and transformed into designed motion representation.
+The interacting hands denoiser further optimizes the interaction process.
+(Right) Overview of Downstream Applications.
+When the control is key frames, we generate the final motions $\mathbf{\hat{x}_{0}^{1:N}}$ by giving the first 5 frames $\mathbf{x_{GT}^{:5}}$ and the last 5 frames $\mathbf{x_{GT}^{-5:}}$ before denoising.
+When the control os wrists’ trajectories, we generate the final motion $\mathbf{\hat{x}_{0}}$ by giving the root angular velocity and linear velocity in each frame before denoising.
+Other controls like 2D keypoints has been illustrated in Appendix due to the space limitation.
+An overview of HandDiffuse is shown in Figure 5.
+With HandDiffuse12.5M dataset, the goal is to synthesize two hands motion $\mathbf{x^{1:N}}$ without or with designed controllers $c$.
+These selected controllers are common in body motion generation and are also able to dictate the synthesis of interacting hands motions, such as key frames (motion in-betweening), trajectory of both wrists (trajectory control) and 2D keypoints(3D hands reconstruction).
+In contrast to single-person human motion generation, the task of generating interacting hands motion requires the network to learn not only the local motion of each hand but also the dynamic process of global interaction.
+Therefore, we adopts a modular design to learn the local pose of single hand and dynamic interacting process separately.
+We also propose two specific hand motion representations for different controllers and Interaction Loss ($\mathbf{Loss}_{interaction}$) to model this dynamic interacting process better.
 Inspired by recent work, we adopt the diffusion model as the backbone which has shown remarkable generative ability in image generation and human motion generation.
@@
 $\beta_{t}$ is the cosine noise variance schedule and $q(\mathbf{x_{t}})$ is near $\mathcal{N}(0,I)$ when t is big enough.
-Then, $\mathbf{x_{t}}$ is sent to a denoiser conditioned on timestep $t$ and different conditions $c$ to predict $\mathbf{\hat{x}_{0}}$:
+Then, $\mathbf{x_{t}}$ is sent to a denoiser $\mathcal{G}$ conditioned on timestep $t$ and different conditions $c$ to predict $\mathbf{\hat{x}_{0}}$:
 $\mathbf{\hat{x}_{0}}=\mathcal{G}(\mathbf{x_{t}},t,c).$
-We use an encoder-only transformer to construct our denoiser $D$ following MDM[51] and we concatenate the left hand and right hand together as input.
-4.1 Motion representation for two hands.
-For different tasks, we have proposed two specific motion representation methods, which we refer to as local representation (local rep) and global representation (global rep).
-The local rep is inspired by HumanML3D[14], a widely used approach in human motion generation.
-In each frame, the pose of left hand is defined by a tuple of $(\dot{R},\dot{T},J_{local},\dot{J}_{local},\theta)$ and the right hand is $(R_{init},T_{init},\dot{R},\dot{T},J_{local},\dot{J}_{local},\theta)$, where $\dot{R}\in\mathbb{R}^{4}$ and $\dot{T}\in\mathbb{R}^{3}$ represent root angular velocity in quaternion and linear velocity respectively.
-$J_{local}\in\mathbb{R}^{3j}$ and $\dot{J}_{local}\in\mathbb{R}^{3j}$ denote the local joint positions and velocities in root space, with $j=21$ represents the number of joints.
-$\theta\in\mathbb{R}^{4i}$ represents the local joint rotations in quaternion representation, following the skeleton structure of MANO with $i=15$.
+Two Denoisers & Interacting Hands DDIM
+HandDiffuse adopts a modular design and applies two denoisers: Single Hand Denoiser(SHDe) and Interacting Hands Denoiser(IHDe) to focus on the different parts of the whole interacting process as shown in Figure 5.
+SHDe is used to learn the characteristics of local hand motion, so we set the global translation and global rotation of the hand to 0, and extract the local joints’ positions ($J_{local},\dot{J}_{local}$) and pose($\theta$) as input.
+By mirroring the right hand to the left hand, we only trained one SHDe.
+By minimizing the
+$\mathbf{Loss}_{reconSH}=\|L_{GT}-\hat{L}_{0}\|_{2},$ (1)
+where $\hat{L}_{0}$ is the denoised single hand parameter and $L_{GT}$ is the ground truth, we model the local motion of single hand.
+After training the SHDe and freezing its parameters, we concatenate the generated local motions of two hands with the global information which is t-step noise and transform the whole vector to designed motion representation.
+IHDe focuses on the global information and will also finetune the local motion.
+During the training stage, after completing a round of denoising, we minimize
+$\mathbf{Loss}_{IHDe}=\mathbf{Loss}_{reconDH}+\mathbf{Loss}_{interaction},$ (2)
+$\mathbf{Loss}_{reconDH}=\|x_{GT}-\hat{x}_{0}\|_{2},$ (3)
+where $\hat{x}_{0}$ is the denoised two hands parameter and $x_{GT}$ is the ground truth, and $\mathbf{Loss}_{interaction}$ will be illustrated later.
+In the inference stage, after denoising, we re-add the noise of $t-1$ steps.
+We refer to this process as Interacting Hands DDIM.
+Subsequent evaluations have proven the excellence of the modular design.
+In terms of model architecture, we adopted an encoder-only transformer similar to MDM(Tevet et al. 2022b). SHDe and IHDe have the same structure but different dimensions.
+Motion Representations for Interacting Hands.
+For different tasks, we have proposed two specific motion representations, which we refer to as local representation (local rep) and global representation (global rep).
+The local rep is inspired by HumanML3D(Guo et al.
+2022), a widely used approach in human motion generation.
+In each frame, the pose of left hand is defined by a tuple of $(\dot{R},\dot{T},J_{local},\dot{J}_{local},\theta)$ and the right hand is $(R_{init},T_{init},\dot{R},\dot{T},J_{local},\dot{J}_{local},\theta)$, where $\dot{R}\in\mathrm{R}^{4}$ and $\dot{T}\in\mathrm{R}^{3}$ represent root angular velocity in quaternion and linear velocity respectively.
+$J_{local}\in\mathrm{R}^{3j}$ and $\dot{J}_{local}\in\mathrm{R}^{3j}$ denote the local joint positions and velocities in root space, with $j=21$ represents the number of joints.
+$\theta\in\mathrm{R}^{4i}$ represents the local joint rotations in quaternion representation, following the skeleton structure of MANO with $i=15$.
 Each sequence is aligned with the left wrist of the first frame as the coordinate system origin.
-Both left and right hands are rotated to align the left-hand R vector with the direction (1, 0, 0) which is shown in Figure 4.
-Therefore, the right hand has $R_{init}\in\mathbb{R}^{4}$ and $T_{init}\in\mathbb{R}^{3}$.
-On the other hand, the global rep is defined as a tuple of ($J_{global}$, $\dot{J}_{global}$, $\theta$) for both hands, where $J_{global}$ represents the positions of hand joints in the world coordinate system and two hands are normalized by employing the same approach as local rep.
-Figure 5: Module of contact potential loss in Interaction Loss.
-The distance between joints has direction.
-Despite the approximate similarity in length of $\overrightarrow{d_{1}}$ and $\overrightarrow{d_{2}}$, their directions are opposite.
-Therefore, it is necessary to consider both the length and direction simultaneously to eliminate ambiguity.
-Figure 6: The Downstream Application of HandDiffuse. The generation results of downstream applications of HandDiffuse. The generated dual-hand motion exhibits natural and realistic hand interaction.
-During the experimental process, we observed that the local rep performed better in the task of trajectory control, where the global rotation and translation of the hand are given.
-The global rep achieved better results in the unconditional generation and motion in-betweening tasks.
-In Section 5.3, a detailed comparison between the two representations is provided.
-4.2 Interaction Loss.
-The first part of the Interaction Loss is the dual-hand joint contact potential loss.
+Both left and right hands are rotated to align the left-hand $R$ vector with the direction (1, 0, 0) which is shown in 5.
+Therefore, the right hand has $R_{init}\in\mathrm{R}^{4}$ and $T_{init}\in\mathrm{R}^{3}$.
+The global rep is defined as a tuple of ($J_{global}$, $\dot{J}_{global}$, $\theta$) for both hands, where $J_{global}$ represents the positions of hand joints in the world coordinate system and two hands are normalized by employing the same approach as local rep.
+During the experimental process, we observed that the local rep performed better in the task where the global rotation and translation of the hand are given.
+The global rep achieved better results in the unconditional generation, the control of key frames(motion in-betweening).
+Figure 6: Two Downstream Applications of HandDiffuse. The generation results of downstream applications of HandDiffuse.
+Interaction Loss   $\mathbf{Loss}_{interaction}$
+The first part of the $\mathbf{Loss}_{interaction}$ is the contact potential loss($\mathbf{Loss}_{potential}$ ).
 In each frame, we first calculate a $25\times 25$ distance matrix $\mathbf{D}$ between the 21 left-hand joints and 21 right-hand joints with 4 more joints sampled in each palm.
 Since the interaction between the hands involves fine-grained movements, we should focus on situations where the finger joints are very close to each other.
-Inspired by the concept of Contact Potential Field[56], we further represent the distances using the elastic potential energy of a spring model and it can be described as
-$\mathbf{P}otential=1/2\mathbf{K}\ relu(\mathbf{\tau}-\|\mathbf{D}_{pred}\|_{2})^{2},$ (1)
+Inspired by the concept of Contact Potential Field(Yang et al.
+2021), we further represent the distances using the elastic potential energy of a spring model and it can be described as
+$\mathbf{P}=1/2\mathbf{K}\ relu(\mathbf{\tau}-\|\mathbf{D}_{pred}\|_{2})^{2},$ (4)
 where $\mathbf{K}$ represents the coefficient of elasticity and $\mathbf{\tau}$ denotes the distance threshold.
@@
 $\mathbf{Loss}_{potential}=|\mathbf{P}_{pred}-\mathbf{P}_{gt}|(1+\overrightarrow{\mathbf{D}_{pred}}\times\overrightarrow{\mathbf{D}_{gt}}),$ (5)
-where (×)($\times$)( × ) means cross product. More is shown in Figure 5.
+where (×)($\times$)( × ) means cross product. More is shown in Appendix.
 Figure 7: Qualitative comparison for different methods.
-For Unconstrained Generation, MDM[52] tend to generate motion with hands shifting away from each other, while InterGen[31] would generate motions with a mean pose.
+For Unconstrained Generation, MDM tend to generate motion with hands shifting away from each other, while InterGen would generate motions with a mean pose.
 In terms of Trajectory Control, MDM yields satisfactory results, whereas InterGen generates motion without incorporating interaction.
-During our testing process, we observed that MANO has strict requirements for the proportional lengths of hand bones.
-Therefore, the second part of the Interaction Loss is the Shape Loss.
+We observed that MANO has strict requirements for the bone length.
+Therefore, the second part of the $\mathbf{Loss}_{interaction}$ is the Shape Loss($\mathbf{Loss}_{shape}$ ).
 We ensure the rationality of hand shape by calculating the differences in bone lengths between the left and right hands, as well as the differences between the predicted shape and the ground truth (GT).
@@
 $\mathbf{Loss}_{shape}=|\mathbf{BL}_{left}-\mathbf{BL}_{right}|+|\mathbf{BL}_{pred}-\mathbf{BL}_{GT}|,$ (6)
-where $BL$ denotes the Euclidean distance of the bones.
-4.3 Downstream Application
-For the task of unconstrained generation and motion in-betweening, the key challenge lies in constructing the global information of the dynamic interaction process between the two hands.
-Therefore, we choose to use the global rep.
-In motion in-betweening task, we provide the ground truth for the previous five frames $\mathbf{x_{0}^{1:5}}$ and the last five frames $\mathbf{x_{0}^{-5:-1}}$ for the diffused motion $\mathbf{x_{t}^{1:N}}$.
+where $BL$ denotes the bone length.
+Downstream Applications
+For the task of key frames control, we provide the ground truth for the previous five frames $\mathbf{x_{GT}^{:5}}$ and the last five frames $\mathbf{x_{GT}^{-5:}}$ for the diffused motion $\mathbf{x_{t}^{1:N}}$.
 For the task of trajectory control, the goal is to generate the reasonable local motion given the global rotation and translation of both hands.
-In this case, using the local rep allows the network to focus on the local motion during the interaction process.
-Furthermore, we provide $\dot{R}^{1:N}$ and $\dot{T}^{1:N}$ for the diffuse $\mathbf{x_{t}^{1:N}}$.
-5 Experiment
-In this section, we conduct several experiments to evaluate our dataset and method for dual-hand motion generation tasks.
-We particularly evaluate (1) the comparison of our method against previous state-of-the-art approaches, (2) the ability of our method to augment existing datasets, and (3) the effectiveness of our hand motion representation and Interaction Loss.
-The above evaluations are conducted specifically on the three tasks mentioned in Section 4.3.
-The models have been trained with $T=1000$ noising steps and a cosine noise schedule.
-All of them have been trained on a single NVIDIA GeForce RTX 3080 GPU for about two days.
-Methods Unconstrained In-betweening Trajectory Control
+In this case, we provide $\dot{R}^{1:N}$ and $\dot{T}^{1:N}$ as control signal for the diffused motion $\mathbf{x_{t}^{1:N}}$.
+The results are presented in Figure 6.
+We also explore HandDiffuse by applying the control of 2D keypoints, and the experiments have shown that HandDiffuse can also solve the task of hands reconstruction.
+Due to the space limitation, the results are presented in Appendix.
+Experiment
+In this section, we conduct several experiments to evaluate our method HandDiffuse for the task of interacting hands motion generation.
+We particularly evaluate (1) the comparison of our method against previous state-of-the-art approaches, and (2) the ablation study.
+The above evaluations are conducted specifically on the two controllers (key frames and trajectory) and other controller(2D keypoints) is evaluated in Appendix due to the space limitation.
+More experiments on the ability of HandDiffuse to augment existing datasets are presented in Appendix.
+The models have been trained with $T=1000$ diffusing steps and a cosine noise schedule.
+The denoiser has 8 layers, 4 heads and the feedforward dimension is 1024.
+We set the length of generated motion $N$ to 200 in all experiments.
+All of them have been trained on a single NVIDIA GeForce RTX 3090 GPU for about two days.
+Method Unconditioned In-betweening Trajectory Control
 […]
-Real 0.050 12.034
-MDM[52] 0.526 10.366 0.616 10.258 0.205 11.265
-InterGen[31] 0.251 4.928 0.269 5.385 0.402 9.102
-Ours 0.176 11.786 0.282 11.533 0.182 11.523
+real 0.050 12.034 1.146
+MDM(Tevet et al. 2023) 0.526 10.366 1.74 0.616 10.258 1.621 0.205 11.265 1.572
+InterGen(Liang et al. 2023) 0.251 4.928 0 0.269 5.385 1.032 0.402 9.102 3.572
+BUDDI(Müller et al. 2024) 0.853 8.282 2.293 1.211 7.341 2.482 0.921 7.497 1.842
+InterHandGen(Lee et al. 2024) 0.437 10.238 1.731 0.734 8.620 1.819 0.539 10.124 1.282
+HandDiffuse(ours) 0.161 11.829 1.553 0.273 11.554 1.562 0.173 11.758 1.141
 Table 2: Quantitative comparisons of different methods.
 We compare the FID and Diversity for different methods.
-’normal-$\downarrow$’ indicates that results are better with lower metrics.
+$\downarrow$’ indicates that results are better with lower metrics, $\uparrow$ is on the contrary.
 Bold means the best result.
-5.1 Comparisons of different methods
-To the best of our knowledge, our proposed method is the first to specifically target hand-only interaction motion generation.
-To provide a comprehensive evaluation, we compare our approach with two recent state-of-the-art human motion generation methods, namely MDM [52] and InterGen [31].
+Task Unconditioned In-betweening Trajectory Control
+[…]
+real 0.050 1.146 - - - -
+w Global rep. 0.161 1.553 0.273 1.562 0.418 1.347
+w Local rep. 0.431 1.962 0.567 1.971 0.173 1.141
+w/o SHDe 0.168 1.575 0.282 1.571 0.182 1.538
+w/o ${Loss}_{potential}$ 0.171 1.606 0.317 1.628 0.210 1.614
+w/o ${Loss}_{shape}$ 0.390 1.583 0.512 1.579 0.380 1.245
+Complete model 0.161 1.553 0.273 1.562 0.173 1.141
+Table 3: Quantitative comparisons. “w/o SHDe” means IHDe is the only denoiser. SDF is only calculated on frames with penetraion.
+Comparisons of different methods
+We compare our approach with MDM (Tevet et al.
+2023)(single human motion generation), InterGen (Liang et al.
+2023)(interacting human motion generation), BUDDI(Müller et al.
+2024)(latent representation of interacting human) and InterHandGen(Liang et al.
+2023)(interacting hands motion generation of single frame).
+The metrics we applied are common used Fréchet Inception Distance (FID), Diversity and sdf loss(SDF).
+FID can measure the realistic of the generated motion.
+Diversity can prove that the motion do not maintain the mean pose for the sake of preserving realistic.
+SDF would take the positive value while zero otherwise, when the hand joints penetrate with each other.
+Using the formulation, we can measure the severity of penetration during the whole motion.
+The strategy we applied to compute these metrics is inspired by MotionCLIP(Tevet et al.
+2022a) and is illustated in Appendix.
 We modify these methods by adapting their motion representation to enable hand motion generation.
+Specifically, we adapted InterHandGen to generate continuous frames following its original pipeline in which we first generated 200 frames of left-hand motion, then generated the right-hand motion based on the left hand.
+We adapted BUDDI for continuous frame generation and treat the motion of each frame as a token, and retained the ”Person Embed” component to distinguish different hands.
 As depicted in Figure 7, both MDM and InterGen exhibit limitations when generating hand motion.
-They tend to produce motion sequences with severe joint penetration or unnatural interaction since they were not originally designed for dual-hand motion generation.
-In contrast, our proposed method showcases superior performance as demonstrated by the lowest Fréchet Inception Distance (FID) values, as shown in Table 2.
-The way we adopted to calculate FID is introduced in the Appendix.
-Figure 8: Cross data Evaluation.
-Motion In-betweening task result on InterHand2.6M[36].
-We sample two frames in InterHand2.6M as the start frame and end frame and generate the in-betweening motion.
-Our HandDiffse can help augment the existing datasets.
-5.2 Cross Datasets Evaluation
-To evaluate our method cross datasets, we conduct experiments on two large-scale dual-hand interaction motion datasets, InterHand2.6M[36] and Re:InterHand [37] by sampling control signal from these two datasets.
-As shown in Figure 8, given the first frame and last frame from InterHand2.6M, our model is able to generate diverse motions.
-Furthermore, Table 3 shows that after performing In-betweening and Trajectory Control tasks on the datasets, the generated data can increase the diversity.
-This demonstrates the effectiveness of our approach and HandDiffuse can contribute to data augmentation for other datasets.
-More results and the details of experiments are introduced in the Appendix.
-Datasets InterHand Re:InterHand
-[…]
-Real 0.113 8.478 0.034 10.583
-In-betweening 0.194 8.171 0.208 11.200
-Root Control 0.142 10.146 0.236 12.979
-Table 3: Cross Dataset Evaluation.The FID and Diversity both increase, indicating that we get a larger dataset from the original dataset by performing data generation tasks.
-5.3 Ablation Study
-As we propose two kinds of hand motion representations in section 4.1, here we demonstrate the effectiveness of the representations in different tasks by conducting qualitative and quantitative comparisons.
-Figure 9 illustrates the limitations of the model without Shape Loss and Interaction Loss.
-By incorporating the Interaction Loss, we mitigate the issue of joint penetration, resulting in improved motion generation quality.
-To provide a quantitative analysis, we employ FID and the penetration metric sdf loss(SDF)[23] for evaluation.
-SDF represents the Signed Distance Field, which could be defined as:
-$\phi(x,y,z)=-\min(SDF(x,y,z),0),$ (4)
-where $(x,y,z)$ denotes the position of MANO vertex.
-According to the definition, when the hand joints penetrate with each other, $\phi$ would take the positive value while zero otherwise.
-Using the formulation, we can measure the severity of penetration during the whole motion.
-Table 4 demonstrates that our model achieves the highest performance when using the proper motion representation and Interaction Loss.
-Tasks Unconstrained In-betweening Trajectory Control
-[…]
-Real 0.050 1.146 - -
-w Global rep. 0.168 1.575 0.282 0.415
-w Local rep. 0.436 1.988 0.572 0.182
-w/o Interaction Loss 0.176 1.652 0.318 0.216
-Complete model 0.168 1.575 0.282 0.182
-Table 4: Quantitative comparisons of different hand motion representations and w/o Interaction Loss.
-SDF indicates the sdf loss for penetration of joints during the whole motion.
-Note that the mean of sdf loss is only calculated on frames with penetraion.
-6 Conclusion
-Figure 9: Ablation Study. In the absence of Shape Loss, the model generates abnormal hand shapes. Similarly, the absence of Interaction Loss leads to hand interaction motion with severe penetration.
+They tend to produce motion sequences with severe joint penetration, unnatural interaction or static mean pose.
+The resulst of HandDiffuse has the highest diversity as shown in Table 2.
+Because the motion generated by InterGen without condition is static, the FID is lower than us and the penetration does not exist.
+Ablation Study
+At first, we demonstrate the effectiveness of the different motion representations in different tasks in Table 3.
+The qualitative result is shown in Appendix.
+By incorporating the $\mathbf{Loss}_{interaction}$, we mitigate the issue of joint penetration, resulting in improved motion generation quality.
+Table 3 demonstrates that our Interacting Hands DDIM with modular deisgn achieves the highest performance when using the proper motion representation and $\mathbf{Loss}_{interaction}$.
+Conclusion
 We present HandDiffuse12.5M, the largest interacting hands dataset with strong interaction and temporal sequences, to open up the research direction of interacting hands motion generation.
-HandDiffuse12.5M consists of various modalities, including synchronous 50-view videos, 3D keypoints and hand poses.
-Based on HandDiffuse12.5M, we propose a strong baseline method HandDiffuse which is the first interacting hands motion generation approach along with two specific controllers: key frames control and trajectory control.
+We propose a strong baseline method HandDiffuse which is the first interacting hands motion generation approach along with some useful controllers: key frames control, trajectory control and 2D keypoints control.
 Our evaluation in multi-stage demonstrates the benefit of our approach againist others and HandDiffuse fills the blank in human motion generation.
-Appendix
-7 Experiment Details
-7.1 Frechet Inception Distance (FID)
-The Frechet Inception Distance (FID) is a metric used to quantify the dissimilarity between the distributions of generated and ground truth data in latent space.
-In our experiments, we employ FID as a crucial indicator to measure the authenticity and plausibility of our generated data.
-And the formula is
-$FID=|\mu_{gt}-\mu_{gen}|^{2}+\text{Tr}(C_{gt}+C_{gen}-2\sqrt{C_{gt}C_{gen}}),$
-where $\mu_{gt}$ and $\mu_{gen}$ are the mean feature vectors of ground truth and generated motions, respectively; $C_{gt}$ and $C_{gen}$ are the covariance feature matrices of real and generated motions, respectively; Tr()($\cdot$)( ⋅ ) denotes the trace of the matrix.
-In the previous works[51], the feature is extracted by a Inception network which is a classifier.
-However, in the interacting hands datasets, it is challenging to label these complex interactive movements, making it difficult to build a classifier.
-Therefore, taking inspiration from MotionCLIP[50], we propose a self-supervised encoder-decoder architecture for the reconstruction network, as depicted in Figure 10.
-More specifically, we concatenate a $z_{tk}$ vector of equal dimensions to the original motion representation $x^{1:T}$.
-This concatenated sequence $x^{1:T+1}$ is then fed into a transformer encoder.
-From the encoder’s output, we select the corresponding position $z_{x}$ as the feature representation for the input motion segment.
-This obtained feature $z_{x}$ serves as the memory for the decoder.
-Additionally, we use a T-length zero vector that undergoes positional encoding as the target for reconstructing the motion in the decoder.
-The reconstruction loss is computed using the L2 norm.
-During experimentation, we observed that using the global representation of the motion improves the convergence of the network.
-Therefore, in subsequent experiments, we convert the motion to a global representation before calculating the motion’s feature.
-Figure 10: Feature Extractor for FID. Inspired by MotionCLIP[50], we train a transformer with encoder-decoder architecture by using self-supervised strategy.
-7.2 Diversity
-Diversity measures the variability in the resulting motion distribution.
-The way we followed to calculate Diversity is as same as MDM[51].
-We randomly generate 1000 samples and select 200 of them for calculating.
-8 HandDiffuse12.5M
-Figure 11 has shown the diversity and complexity of our dataset HandDiffuse12.5M. Figure 12 has shown our manual annotation tool.
-Figure 11: Overview of HandDiffuse12.5M. Each image represents one frame of hand motion from a different sequence.
-Figure 12: Annotation Tool. Our specially designed tool for hand joint annotation of severely occluded frames.
-9 More Results
-Figure 13 has shown the results of trajectory control task.
-In each sequence, we adopt the wrists’ rotation and translation from InterHand2.6M[36] as our control signal and we can generate different motion based on this trajectory.
-Figure 13: Results of Trajectory Control by Sampling from InterHand2.6M. Note that only Root orientation and position is given, the local motions for hand joints are generated from our model.
+Acknowledgments
+We would like to express our sincere gratitude to all the students and faculty members of the VDI Laboratory at ShanghaiTech University for their invaluable assistance and guidance throughout this experiment.
+We also appreciate the use of the filming equipment provided by the VDI Laboratory.
+Additionally, we would like to acknowledge our collaborators: Sihang Xu, Hongdi Yang, Yiran Liu, Xin Chen, Jingya Wang, Jingyi Yu, and Lan Xu, for their contributions to this work.